# Unleashing the Power of CNN and Transformer for Balanced RGB-Event Video Recognition

### v1

@@ IV Experiments
IV-A Datasets and Evaluation Metric
@@
In our experiments, two large-scale frame-event based pattern recognition datasets are used, including HARDVS [5] and PokerEvent [12]. Specifically speaking, the HARDVS  [^3] https://github.com/Event-AHU/HARDVS dataset is collected using a DVS346 event camera, which concentrates on recognizing human activities such as walking, running, and crouching. It consists of 300 classes and includes 96,908 RGB-Event samples. It is divided into training and testing subsets, with 64,522 and 32,386 samples, respectively. The PokerEvent  [^4] https://github.com/Event-AHU/SSTFormer dataset focuses on recognizing character patterns in poker cards. It consists of 114 classes and includes 24,415 RGB-Event samples recorded using a DVS346 event camera. The dataset is divided into training and testing subsets, with 16,216 and 8,199 samples, respectively.
The top-1 accuracy and top-5 accuracy are adopted to evaluate the performance of our model and other state-of-the-art recognition algorithms.
IV-B Implementation Details
@@
In our experiments, we set the batch size to 8 and train our framework for 30 epochs on both the HARDVS and PokerEvent datasets. We use the SGD [46] optimizer with an initial learning rate of 0.001 for the training process. Our parameter scheduler is MultiStepLR, and when the training epoch reaches 25, the learning rate decreases by a factor of 0.1. During the training phase, we first perform MultiScaleCrop on the input images of two modalities, with scales set to 1, 0.875, 0.75, and 0.66. Then, we resize them to $224\times 224$, and finally apply a flip with ratio of 0.5. We use RGB and Event videos for the two branches in the framework, with 8 frames selected for each modality, totaling 16 frames. We apply a temporal shift operation to the input feature, with the number of divisions for the shift set to 8. We select the widely used cross-entropy loss as the loss function to quantify the disparity between the ground truth and our model’s predictions. This loss function is commonly employed in classification tasks and is effective in penalizing the deviations of predicted probabilities from the actual class labels. By utilizing the cross-entropy loss, we aimed to optimize parameters of our model to minimize the discrepancy between the predicted and actual outcomes, thereby enhancing the predictive performance. Our code is implemented using PyTorch [47] based on Python. All the experiments were conducted on a server with RTX3090 GPUs.
TABLE I: Results on the HARDVS dataset (Event+RGB). The best and second results are highlighted in red and blue.
No. Algorithm Publish Backbone Top-1 |||| Top-5
#01 C3D [48] ICCV-2015 3D-CNN 50.88 |||| 56.51
#02 TSM [14] ICCV-2019 ResNet-50 52.58 |||| 62.12
#03 TimeSformer [49] ICML-2021 ViT 51.57 |||| 58.48
#04 X3D [50] CVPR-2020 ResNet 47.38 |||| 51.42
#05 ESTF [5] AAAI-2024 ResNet18 49.93 |||| 55.77
#06 SSTFormer [12] arXiv-2023 SNN-Former 52.97 |||| 60.17
#07 TSCFormer (Ours) - TSCFormer 53.04 |||| 62.67
IV-C Comparison on Public Benchmarks
@@
To objectively evaluate the performance of our proposed method, we conduct extensive experiments on HARDVS and PokerEvent datasets and compare our results with other state-of-the-art recognition models in Table I and Table II.
@@
Results on HARDVS Dataset [5] As shown in Table I, we compare our proposed TSCFormer with 6 SOTA models, including 3D CNN based model C3D [48], ResNet based model (TSM [14], X3D [50] and ESTF [5]), Transformer-based methods (TimeSformer [49] and SSTFormer [12]). It is easy to find that our TSCFormer achieves the best recognition performance on this dataset, i.e., $53.04\%$ on the top-1 evaluation metric. For the top-5 metric, our model also achieves the best accuracy, i.e., $62.67\%$, which exceeds the second one (TSM [14]) by a margin $+0.55\%$. These results fully validated the effectiveness of our proposed local-global feature aggregation mechanism for RGB-Event pattern recognition.
TABLE II: Results on the PokerEvent dataset (Event+RGB). The best and second results are highlighted in red and blue.
No. Algorithm Publish Backbone Top-1
#01 C3D [48] ICCV-2015 3D-CNN 51.76
#02 TSM [14] ICCV-2019 ResNet-50 55.43
#03 ACTION-Net [51] CVPR-2021 ResNet-50 54.29
#04 TAM [15] ICCV-2021 ResNet-50 53.65
#05 V-SwinTrans [16] CVPR-2022 Swin Transformer 54.17
#06 TimeSformer [49] ICML-2021 ViT 55.69
#07 X3D [50] CVPR-2020 ResNet 51.75
#08 MVIT [52] CVPR-2022 ViT 55.02
#09 SSTFormer [12] arXiv-2023 SNN-Former 54.74
#10 TSCFormer (Ours) - TSCFormer 57.70
@@
Results on PokerEvent Dataset [12] As shown in Table II, we compare our model with 9 SOTA algorithms on the PokerEvent dataset. We can find that our model achieves $57.7\%$ on the top-1 evaluation metric which surpasses the latest recognition methods V-SwinTrans [16] and MVIT [52]. Also, our model beats the baseline TSM [14] which obtained $55.43\%$ on this dataset. The detailed comparisons of recognition results of five classes (White bear, Ostrich, Cyclophiops, Chameleon, Alligator) on PokerEvent dataset are illustrated in Fig. 7. We can find that our model significantly improves the recognition results over the TSM model on these 5 categories. To be specific, the top-1 accuracy are improved up to $97.87\%$, $98.00\%$, $98.00\%$, $97.87\%$, $95.92\%$, respectively. These experimental results further proved the effectiveness of our proposed local-global feature fusion and enhancement modules for RGB-Event based video recognition.
@@ IV-D Ablation Study
TABLE III: Component Analysis on the PokerEvent dataset.
@@
Component Analysis.  The key components of our proposed TSCFormer framework are the Temporal Shift CNN Sub-Network (T.S.), Cross-attention based Bridge Module, and Self-attention based Transformer layer. As shown in Table III, we conduct the component analysis on the PokerEvent dataset by removing each module separately. Specifically, we can find that our model achieves $57.70\%$ on the top-1 accuracy when all these modules are used. The results are dropped to $56.71\%$, $56.06\%$, and $56.49\%$ when removing the temporal shift module, Transformer, and Bridge module, respectively. These results demonstrate that all the modules contribute to our final recognition performance. Also, we find that when only T.S. is used, the result is $56.88\%$, however, the result is dropped when introducing the Transformer layers only. If we use the Transformer and Bridge module simultaneously, the results can be improved to $57.70\%$, which demonstrates that the cross-attention based Bridge module is a vital component in the TSCFormer. When only the CNN is used for recognition, the top-1 accuracy is $55.93\%$ which validated the effectiveness of the temporal shift module in our framework.
TABLE IV: Results of different inputs on PokerEvent dataset.
Input Backbone Top1
@@
Influence of Single- and Multi-modal Inputs.  To validate the effect of different modalities and the combination of RGB and Event data on recognition performance, we conduct experiments using various input combinations. As shown in Table IV, using only event data achieved the lowest accuracy, indirectly indicating the difficulty of extracting features from Event data compared to RGB data. Additionally, using only RGB data achieved an accuracy of $56.64\%$, while combining both modalities yielded the highest result. This suggests the beneficial impact of multi-modal data on improving recognition performance and the effectiveness of our proposed method in thoroughly extracting and integrating features from two modalities.
@@ IV-E Parameter Analysis
IV-F Visualization
@@
Feature Maps and Distribution.  As shown in Fig. 10, we give a comparison between the feature maps of ResNet50, TSM, and our TSCFormer. We can find that our model achieves more focused attention on the Poker samples. For the feature maps obtained on the HARDVS dataset, as shown in Fig. 11, we can find that our model also can locate the key regions that will help recognize human activities. As illustrated in Fig. 12, we visualize the feature distribution of TSM and our TSCFormer on the PokerEvent dataset. We showcase the distances between different classes (randomly 10 classes are selected). It is evident that the feature aggregation effect of our TSCFormer is more precise than that of TSM, effectively minimizing the distribution error observed in the baseline. This further underscores the superiority of our method over the baseline approach.
@@
IV-G Limitation Analysis
Although our proposed TSCFormer achieves better results on both HARDVS and PokerEvent datasets, however, our framework can still be improved from the following aspects: Firstly, our TSCFormer contains the ResNet as the backbone network and is initialized using the weights pre-trained on the ImageNet dataset. However, it failed to reap the benefits of pre-trained large models, and a lot of work has demonstrated its strong generalization ability [31, 18]. Secondly, we formulate the RGB-Event pattern recognition problem as a mapping from vision modality to semantic labels in this work, however, there still exists the semantic gaps between these modalities. For example, the BLIP-2 [53] demonstrates that the introduction of Q-Former will bridge such semantic gaps significantly. In our future works, we will consider further improving our framework from the aforementioned two aspects.
V Conclusion and Future Works

### v2

@@ 4 Experiments
4.1 Datasets and Evaluation Metric
@@
In our experiments, two large-scale frame-event based pattern recognition datasets are used, including HARDVS wang2022hardvs and PokerEvent wang2023sstformer . Specifically speaking, the HARDVS  [^3] https://github.com/Event-AHU/HARDVS dataset is collected using a DVS346 event camera, which concentrates on recognizing human activities such as walking, running, and crouching. It consists of 300 classes and includes 96,908 RGB-Event samples. It is divided into training and testing subsets, with 64,522 and 32,386 samples, respectively. The PokerEvent  [^4] https://github.com/Event-AHU/SSTFormer dataset focuses on recognizing character patterns in poker cards. It consists of 114 classes and includes 24,415 RGB-Event samples recorded using a DVS346 event camera. The dataset is divided into training and testing subsets, with 16,216 and 8,199 samples, respectively. The top-1 accuracy and top-5 accuracy are adopted to evaluate the performance of our model and other state-of-the-art recognition algorithms.
4.2 Implementation Details
@@
In our experiments, we set the batch size to 8 and train our framework for 30 epochs on both the HARDVS and PokerEvent datasets. We use the SGD Bottou2012Stochastic optimizer with an initial learning rate of 0.001 for the training process. Our parameter scheduler is MultiStepLR, and when the training epoch reaches 25, the learning rate decreases by a factor of 0.1. During the training phase, we first perform MultiScaleCrop on the input images of two modalities, with scales set to 1, 0.875, 0.75, and 0.66. Then, we resize them to $224\times 224$, and finally apply a flip with ratio of 0.5. We use RGB and Event videos for the two branches in the framework, with 8 frames selected for each modality, totaling 16 frames. We apply a temporal shift operation to the input feature, with the number of divisions for the shift set to 8. We select the widely used cross-entropy loss as the loss function to quantify the disparity between the ground truth and our model’s predictions. This loss function is commonly employed in classification tasks and is effective in penalizing the deviations of predicted probabilities from the actual class labels. By utilizing the cross-entropy loss, we aimed to optimize parameters of our model to minimize the discrepancy between the predicted and actual outcomes, thereby enhancing the predictive performance. Our code is implemented using PyTorch paszke2019pytorch based on Python. All the experiments were conducted on a server with RTX3090 GPUs.
4.3 Comparison on Public Benchmarks
@@
To objectively evaluate the performance of our proposed method, we conduct extensive experiments on HARDVS and PokerEvent datasets and compare our results with other state-of-the-art recognition models in Table 1 and Table 2.
@@
Results on HARDVS Dataset wang2022hardvs As shown in Table 1, we compare our proposed TSCFormer with 6 SOTA models, including 3D CNN based model C3D tran2015learning , ResNet based model (TSM lin2019tsm , X3D feichtenhofer2020x3d and ESTF wang2022hardvs ), Transformer-based methods (TimeSformer bertasius2021space and SSTFormer wang2023sstformer ). It is easy to find that our TSCFormer achieves the best recognition performance on this dataset, i.e., $53.04\%$ on the top-1 evaluation metric. For the top-5 metric, our model also achieves the best accuracy, i.e., $62.67\%$, which exceeds the second one (TSM lin2019tsm ) by a margin $+0.55\%$. These results fully validated the effectiveness of our proposed local-global feature aggregation mechanism for RGB-Event pattern recognition.
@@
Results on PokerEvent Dataset wang2023sstformer As shown in Table 2, we compare our model with 9 SOTA algorithms on the PokerEvent dataset. We can find that our model achieves $57.7\%$ on the top-1 evaluation metric which surpasses the latest recognition methods V-SwinTrans liu2022video and MVIT li2022mvitv2 . Also, our model beats the baseline TSM lin2019tsm which obtained $55.43\%$ on this dataset. The detailed comparisons of recognition results of five classes (White bear, Ostrich, Cyclophiops, Chameleon, Alligator) on PokerEvent dataset are illustrated in Fig. 6. We can find that our model significantly improves the recognition results over the TSM model on these 5 categories. To be specific, the top-1 accuracy are improved up to $97.87\%$, $98.00\%$, $98.00\%$, $97.87\%$, $95.92\%$, respectively. These experimental results further proved the effectiveness of our proposed local-global feature fusion and enhancement modules for RGB-Event based video recognition.
Table 1: Results on the HARDVS dataset (Event+RGB). The best and second results are highlighted in red and blue.
No. Algorithm Publish Backbone Modality Top-1 Top-5
#01 C3D tran2015learning ICCV-2015 3D-CNN Event+RGB 50.88 56.51
#02 TSM lin2019tsm ICCV-2019 ResNet-50 Event+RGB 52.58 62.12
#03 TimeSformer bertasius2021space ICML-2021 ViT Event+RGB 51.57 58.48
#04 X3D feichtenhofer2020x3d CVPR-2020 ResNet Event+RGB 47.38 51.42
#05 ESTF wang2022hardvs AAAI-2024 ResNet18 Event+RGB 49.93 55.77
#06 SSTFormer wang2023sstformer arXiv-2023 SNN-Former Event+RGB 52.97 60.17
#07 TSCFormer (Ours) - TSCFormer Event+RGB 53.04 62.67
Table 2: Results on the PokerEvent dataset (Event+RGB). The best and second results are highlighted in red and blue.
No. Algorithm Publish Backbone Modality Top-1
#01 C3D tran2015learning ICCV-2015 3D-CNN Event+RGB 51.76
#02 TSM lin2019tsm ICCV-2019 ResNet-50 Event+RGB 55.43
#03 ACTION-Net wang2021action CVPR-2021 ResNet-50 Event+RGB 54.29
#04 TAM liu2021tam ICCV-2021 ResNet-50 Event+RGB 53.65
#05 V-SwinTrans liu2022video CVPR-2022 Swin Transformer Event+RGB 54.17
#06 TimeSformer bertasius2021space ICML-2021 ViT Event+RGB 55.69
#07 X3D feichtenhofer2020x3d CVPR-2020 ResNet Event+RGB 51.75
#08 MVIT li2022mvitv2 CVPR-2022 ViT Event+RGB 55.02
#09 SSTFormer wang2023sstformer arXiv-2023 SNN-Former Event+RGB 54.74
#10 TSCFormer (Ours) - TSCFormer Event+RGB 57.70
@@ 4.4 Ablation Study
Component Analysis.  The key components of our proposed TSCFormer framework are the Temporal Shift CNN Sub-Network (T.S.), Cross-attention based Bridge Module, and Self-attention based Transformer layer. As shown in Table 3, we conduct the component analysis on the PokerEvent dataset by removing each module separately. Specifically, we can find that our model achieves $57.70\%$ on the top-1 accuracy when all these modules are used. The results are dropped to $56.71\%$, $56.06\%$, and $56.49\%$ when removing the temporal shift module, Transformer, and Bridge module, respectively. These results demonstrate that all the modules contribute to our final recognition performance. Also, we find that when only T.S. is used, the result is $56.88\%$, however, the result is dropped when introducing the Transformer layers only. If we use the Transformer and Bridge module simultaneously, the results can be improved to $57.70\%$, which demonstrates that the cross-attention based Bridge module is a vital component in the TSCFormer. When only the CNN is used for recognition, the top-1 accuracy is $55.93\%$ which validated the effectiveness of the temporal shift module in our framework. We analyze the impact on the classification performance of removing the F2V and F2E modules. As shown in Table 4, when the F2V and F2E modules are used, the top-1 accuracy of the model is 57.70. However, when the F2V and F2E modules are removed, the top-1 accuracy of the model drops to 56.82. The experimental results demonstrate the effectiveness of the proposed F2V and F2E modules.
Compare of Only Event Inputs.  Table 5 shows the experimental results of these video recognition approaches when only event data is utilized. Our model achieves better performance compared to other video recognition approaches. The experimental results demonstrate the effectiveness of the proposed method.
Influence of Single- and Multi-modal Inputs.  To validate the effect of different modalities and the combination of RGB and Event data on recognition performance, we conduct experiments using various input combinations. As shown in Table 6, using only event data achieved the lowest accuracy, indirectly indicating the difficulty of extracting features from Event data compared to RGB data. Additionally, using only RGB data achieved an accuracy of $56.64\%$, while combining both modalities yielded the highest result. This suggests the beneficial impact of multi-modal data on improving recognition performance and the effectiveness of our proposed method in thoroughly extracting and integrating features from two modalities.
Table 3: Component Analysis on the PokerEvent dataset.
@@
Table 4: Results of the model with and without the F2V and F2E modules.
No. Method Top-1
01  with F2V and F2E  57.70
02  without F2V and F2E  56.82
Table 5: Results of utilize only event data on the HARDVS dataset.
No. Algorithm Modality Top-1
01  C3D  Event  50.52
02  TSM  Event  52.63
03  TimeSformer  Event  50.77
04  X3D  Event  45.82
05  ESTF  Event  51.22
06  SSTFormer  Event  51.87
07 TSCFormer(Ours)  Event 52.70
Table 6: Results of different inputs on PokerEvent dataset.
Input Backbone Top-1
@@ 4.5 Parameter Analysis
Figure 11: The visual comparison examples of static scenes, high-motion scenarios and high dynamic range environments.
Figure 14: The top-5 prediction examples of static scenes, high-motion scenarios and high dynamic range environments.
4.6 Visualization
@@
Feature Maps and Distribution.  As shown in Fig. 9, we give a comparison between the feature maps of ResNet50, TSM, and our TSCFormer. We can find that our model achieves more focused attention on the Poker samples. For the feature maps obtained on the HARDVS dataset, as shown in Fig. 10, we can find that our model also can locate the key regions that will help recognize human activities.
As shown in the Fig. 11, the first and second rows represent the RGB and event samples and their corresponding visualization effects under static scenes. The third and fourth rows represent the situations under high-motion scenarios. The fifth to eighth rows display the cases in a high dynamic range environment. Among them, the fifth and sixth rows indicate that the samples are under high illumination, and the seventh and eighth rows indicate that the samples are under low illumination. In multiple challenging scenarios of Fig. 11, the sample visualization effects focus on the people in the scenes and their behaviors and actions, indicating that it is effective to perform video classification by combining RGB and event data.
As illustrated in Fig. 12, we visualize the feature distribution of TSM and our TSCFormer on the PokerEvent dataset. We showcase the distances between different classes (randomly 10 classes are selected). It is evident that the feature aggregation effect of our TSCFormer is more precise than that of TSM, effectively minimizing the distribution error observed in the baseline. This further underscores the superiority of our method over the baseline approach.
@@
As shown in Fig. 14, the samples in the first row represent the situation under static scenes. The model predicts the top five classification categories, and the top-1 prediction result is consistent with the ground truth. The samples in the second row are under high-motion scenarios. The samples in the third and fourth rows are in a high dynamic range environment, representing the environments with high illumination and low illumination respectively. In multiple scenarios, the top five prediction results of the model are consistent with the ground truth, which proves that it is beneficial to perform video classification by combining RGB and event data.
4.7 Limitation Analysis
Although our proposed TSCFormer achieves better results on both HARDVS and PokerEvent datasets, however, our framework can still be improved from the following aspects: Firstly, our TSCFormer contains the ResNet as the backbone network and is initialized using the weights pre-trained on the ImageNet dataset. However, it failed to reap the benefits of pre-trained large models, and a lot of work has demonstrated its strong generalization ability radford2021learning ; wang2023MMPTMs . Secondly, we formulate the RGB-Event pattern recognition problem as a mapping from vision modality to semantic labels in this work, however, there still exists the semantic gaps between these modalities. For example, the BLIP-2 li2023blip2 demonstrates that the introduction of Q-Former will bridge such semantic gaps significantly. In our future works, we will consider further improving our framework from the aforementioned two aspects.
5 Conclusion and Future Works
@@
Acknowledgments
This work is supported by the National Natural Science Foundation of China (No. 62102205, 62076004, 62102207, 62332002, 62027804, 62088102); Anhui Provincial Key Research and Development Program under Grant 2022i01020014; The authors acknowledge the High-performance Computing Platform of Anhui University for providing computing resources.
Publish Information
Name of the Journal: Machine Intelligence Research